OpenStudy (anonymous):

A city's population, P (in thousands), can be modeled by the equation P = 130 (1.03)^x, where x is the number of years after January 1, 2000. For what value of x does the model predict that the population of the city will be the approximately 17,000. a. 8 years b. 9 years c. 10 years d. 11 years

OpenStudy (anonymous):

log17000 = log130 +xlog1.03

OpenStudy (anonymous):

So what's the answer or I don't get it..

OpenStudy (anonymous):

?

OpenStudy (anonymous):

start with \[1700=130(1.03)^x\] \[\frac{170}{13}=(1.03)^x\] \[x=\frac{\ln(\frac{170}{13})}{\ln(1.03)}\]

OpenStudy (anonymous):

now you need a calculator

OpenStudy (anonymous):

i get 86 so i assume there is a typo in the question. is it \[P=130(1.03)^x\] ir \[P=1300(1.03)^x\]?

OpenStudy (anonymous):

if it is the second one answer is 9 years

OpenStudy (anonymous):

17000 is the population not 1700

OpenStudy (anonymous):

Its 170,000. Sorry typo ....So it's 9 years?

OpenStudy (anonymous):

it says P is in measured in thousands so 170000 is collapsed to 170. P(thousands)=130*(1.03)^x (170/130)=(1.03)^x log(170/130)=x*log(1.03) x= [log(170/130)/log(1.03)] x=9.0756 years